OpenStudy (narissa):

PLEASE HELP!! The length (L) of a photograph is 8mm longer than its width (W). Its perimeter (P) is more than 60mm. 𝐏 = 𝟐𝐖 + 𝟐𝐋 a) Write an equation for the length in terms of the width L = _____________________ b) Use these expressions to write an inequality based on the given information. c) Solve the inequality clearly indicating the length and width of the photograph. d) Write your answer in a complete sentence

OpenStudy (narissa):

@satellite73

OpenStudy (narissa):

@nincompoop

OpenStudy (narissa):

@Rushwr

OpenStudy (narissa):

@ChantySquirrel1129**

OpenStudy (compassionate):

What is the inequality

OpenStudy (narissa):

i have no idea what a inequality is :/

OpenStudy (narissa):

my brother help me solve (a) but in still confused about (b) (c) and (d)

OpenStudy (compassionate):

Show me your equation

OpenStudy (narissa):

for a?

OpenStudy (narissa):

L = (P - 2W) / 2

OpenStudy (compassionate):

The length (L) of a photograph is 8mm longer than its width (W). Its perimeter (P) is more than 60mm. Length is 8mm mre than it's width So L = w + 8 It's perimeter is 60mm or more. \[2(w + 8) + 2w /\ge 60\]

jhonyy9 (jhonyy9):

@Compassionate - i think more than 60 mm mean strictly greater than 60 and not can being equal 60 - how you think it ?

OpenStudy (narissa):

L=w+8 would be a?

OpenStudy (compassionate):

My logic being, if you take all the sides of a triangle, it will either be equal to or greater than the perimeter. But, hey. 2(w + 8) 2w > 60 will yield the same solution.

jhonyy9 (jhonyy9):

pardon - just for ,,asker" understand it sure clearly - ty !!!

jhonyy9 (jhonyy9):

@nincompoop how you think than the perimeter need being strictly greater than 60 this mean that the length and width of sides need being strictly greater not smaller - how you have wrote above - check it please prove than l < 19 --- 2l < 38 and w<11 --- 2w < 22 ----------------------------------------------- perimeter resulte strictly < than 38+22=60 but in the text of exercise have wrote strictly greater how you think it so ????

OpenStudy (nincompoop):

width is greater than 11 length is width + 8

OpenStudy (nincompoop):

let me re-write so it is not confusing. The length (L) of a photograph is 8mm longer than its width (W). Its perimeter (P) is more than 60mm. 𝐏 = 𝟐𝐖 + 𝟐𝐋 a) Write an equation for the length in terms of the width L = _____________________ b) Use these expressions to write an inequality based on the given information. c) Solve the inequality clearly indicating the length and width of the photograph. d) Write your answer in a complete sentence

OpenStudy (nincompoop):

IF P = 2W + 2L then, 2W + 2L > 60 is the same way as P > 60 in words: perimeter is greater than 60 L = W+8 (a) length in terms of width L-8 = W (*) width 2(W+8) + 2W >60 2W+16 + 2W > 60 4W + 16 > 60 4W > 60 -16 4W > 44 W > 11 W+8 = L 11<W+8 = L Width is greater than 11 Length is W+8

OpenStudy (narissa):

L=W+8?

OpenStudy (nincompoop):

more specifically 11<W+8 = L because there is a restriction in the value of width

OpenStudy (nincompoop):

the value of length is dependent on the value of width the value of width is anything greater than 11 so, we can say that the value of length is greater than 19

OpenStudy (narissa):

okay